OpenStudy (anonymous):

which of the following are solutions to the equation below? Check all that apply. (2x+3)^2=10

OpenStudy (anonymous):

a.\[x=-\sqrt{10}+\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

B.\[x=\frac{ -\sqrt{10-3} }{ 2 } \]

OpenStudy (anonymous):

C.\[x=\frac{ \sqrt{10-3} }{ 2 }\]

OpenStudy (anonymous):

D.\[x=-\frac{ \sqrt{13} }{ 2 }\]

OpenStudy (anonymous):

E.\[x=\frac{ \sqrt{7} }{ 2 }\]

OpenStudy (anonymous):

f.\[x=\sqrt{10}-\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

A and F

OpenStudy (anonymous):

A AND F

OpenStudy (anonymous):

yo still there ilfy214

OpenStudy (anonymous):

I honestly don't know... I'm trying to work it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how did you get A and F?

OpenStudy (anonymous):

so not F, right?

OpenStudy (anonymous):

√(2x + 3)² = ±√10 2x + 3 = ±√10 Finally... x = (-3 ± √10)/2 [A and F]

OpenStudy (anonymous):

so what do you think

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

im guessing.. you know. lol

OpenStudy (anonymous):

Because I got \[x = -3/2 \pm \sqrt(5/2)\]

OpenStudy (radar):